OpenStudy (anonymous):

Use properties of logarithms with the given approximations to evaluate the expression. loga3=.477 and loga2=.301. Use one or both of these values to evaluate loga27

OpenStudy (anonymous):

\[\log_a(27)=\log_a(3^3)=3\times \log_a(3)\]

OpenStudy (anonymous):

so \(3\times .447\) is your answer

OpenStudy (anonymous):

1.341 is the wrong answer so it says

OpenStudy (anonymous):

whatever "it" is, "it" is mistaken

OpenStudy (anonymous):

the website said the answer is 1.431 how is that so?

OpenStudy (anonymous):

because i made a typo. it is \(\log_a(3)=.477\) and \(3\times .477=1.431\)